# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v1

@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
    $\sqrt{n}(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}(0,2(\sigma_{0}^{2}(a_{0})/a_{0}^{1+2\kappa})^{2})$.
We conclude this section with some comments on the reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. In particular, assuming $\mu$ unknown and using (27), for the latter model the associated microergodic parameter is given by $\sigma^{2}B(\kappa,\mu,d)/(\alpha\mu)^{2\kappa+1},$ and for the former model, it is given by $\sigma^{2}/\alpha^{2\kappa+1}.$ Notably, the microergodic parameter in the second case does not depend on $\mu$. This implies that, for the $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ model, we have two microergodic parameters, $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$, that can be estimated consistently under fixed-domain asymptotics. As a consequence, from an inferential point of view, the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ should be preferred over the $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$ model. However, particularly for large datasets, the $\mu$ parameter should be fixed, as it allows the construction of sparse covariance matrices. In this case, the microergodic parameter is given by $\sigma^{2}/\alpha^{2\kappa+1}$ for both parameterizations.
@@ 7 A simulation study
In this section, we perform a simulation study on the estimation of the $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ family. We simulate 500 realizations, using Cholesky decomposition, of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $d=2$, setting $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. We have considered a regular grid with increments $0.025$ over $[0,1]^{2}$, that is $1681$ grid points. Then the grid points have been perturbed, adding a uniform random value on $[-0.01,0.01]$ to each coordinate. We assume the parameters $\kappa$ and $\mu$ known and we focus on the ML estimation of $\sigma^{2}$ and $a$ and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing and fixed domain scenarios. Since the increasing as well as the fixed-domain frameworks can be mimicked by fixing the number of location sites over a given spatial domain and decreasing or increasing the spatial dependence (Zhang and Zimmerman,, 2005; Kaufman and Shaby,, 2013), we set the compact support $a=0.1,0.25,0.5$. We consider $a=0.1$ as an increasing domain scenario and $a=0.5$ as a fixed domain scenario.
Table 1 reports biases and standard deviations (in parentheses) of the 500 ML parameter estimates, namely $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$, $\kappa=0,1,2$, $\mu=4,6,8$ and $a=0.1,0.25,0.5$.
@@
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ $-0.0150$ $-0.0219$ $-0.0151$ $-3.9684$ $-4.4120$ $-3.7659$ $-496.06$ $-486.01$ $-346.94$
@@
Overall, the ML estimation of $\sigma^{2}$ and $a$ is approximately unbiased, and the standard deviations increase when switching from the increasing domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$), as expected. Another general pattern that can be observed is that, for each $\kappa$, increasing the shape parameter $\mu$ results in a lower standard deviation.
For the microergodic parameter, it can be observed that under the increasing domain scenario ($a=0.1$), the ML estimation can be significantly unbiased but with a high standard deviation, particularly when $\kappa$ increases. When moving to the fixed-domain scenario ($a=0.5$), both the bias and the standard deviation are significantly reduced, as expected. In addition, the standard deviation of the microergodic parameter increases as the smoothness parameter $\kappa$ increases.
Figure 6 depicts the boxplots of the ML estimates of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (on the left) and $a=0.5$ (on the right) when $\kappa=0$. The Gaussian distribution of the estimates for $\sigma^{2}$ and $a$ is more evident under the increasing domain scenario, as expected. On the other hand, when switching to the fixed-domain scenario, the distribution of the ML estimates of the microergodic parameter becomes more centered around the true values, and the distribution appears Gaussian, as expected.
@@ 8 A real data application
We consider data from WorldClim (www.worldclim.org), a global database of high spatial resolution weather and climate data for the years 1970–2000, which can be downloaded using the R package geodata (Hijmans et al.,, 2024). Specifically, we analyze mean temperature data for September observed at 725,330 georeferenced locations across Madagascar. To make ML estimation feasible, we select a random sample of 3,500 locations and perform the analysis using a UTM projection.
Following Li and Zhang, (2011), we first detrend the data using splines to remove the cyclic patterns of both variables along the two coordinates. We then treat the residuals $y(\boldsymbol{s}_{i})$, $i=1,\ldots,3500$, as realizations from a zero-mean RF $\{Y(\boldsymbol{s}):\boldsymbol{s}\in D\}$. Figure 7, from left to right, depicts the colored map, the boxplot, and the empirical semivariogram of the residuals. The graphics suggest that a weakly stationary RF with a marginal distribution capable of modeling heavy tails is potentially an appropriate model for our data.
Refer to caption Refer to caption Refer to caption
Figure 7: From left to right: spatial map of the residuals for the Madagascar mean temperature data, boxplot, and empirical semivariogram of the data.
Let $\{Z(\boldsymbol{s}):\boldsymbol{s}\in D\}$ be a standard Gaussian RF with an isotropic parametric correlation function. In our analysis, we consider two specifications for $Y(\boldsymbol{s})$:
  • a zero-mean Gaussian RF, i.e., $Y(\boldsymbol{s})=\sigma Z(\boldsymbol{s})$, and
[…]
@@
[…]
$Time$ $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
[…]
@@
[…]
@@ 9 Concluding Remarks
In this paper, we study the Gauss hypergeometric covariance model, a highly flexible kernel introduced by Emery and Alegría, (2022), which generalizes both the Matérn and Generalized Wendland models. We establish necessary and sufficient conditions for the validity of the model and highlight that the model suffers from identifiability issues under both increasing and fixed domain asymptotics. To address the identifiability problem, we introduce a new class of compactly supported kernels that satisfy a specific optimality criterion. In particular, the proposed model achieves the maximum integral range within the class of Gauss hypergeometric correlations. From this perspective, the proposed model can be regarded as an enhancement of the Generalized Wendland model.
The proposed model has the potential to outperform the Matérn family in terms of model fitting, prediction performance, and computational efficiency, as demonstrated in the real-data application.
Additionally, the proposed model can be easily implemented, as efficient numerical computation of the Gaussian hypergeometric function is available in various libraries, such as the GNU Scientific Library (Gough,, 2009), and in major statistical software packages, including R, MATLAB, and Python. In particular, the R package GeoModels (Bevilacqua et al.,, 2025), used in this paper for numerical experiments and applications, computes the proposed model using the Python implementation of the Gaussian hypergeometric function in the SciPy library (Virtanen et al.,, 2020).
From an inferential perspective, the results on the asymptotic properties of the maximum likelihood estimator of the microergodic parameter could be generalized to include a nugget effect, as in Tang et al., (2021). Additionally, the estimation of the smoothness parameter could, in principle, be studied following the ideas in Loh and Sun, (2023). In a Bayesian setup, the fixed-domain asymptotic posterior distribution of the microergodic parameter could be investigated following Li, (2022). Finally, the proposed model can be generalized in various directions, such as space-time, multivariate, or non-stationary contexts. These topics are left for future research.
@@ Acknowledgement
We would like to thank Professor Igor Kondrashuk for his help in partially proving Lemma 1. This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
@@ A.1 Proof of Theorem 1
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1). Using results in Skorokhod and Yadrenko, (1973) and Ibragimov and Rozanov, (1978), Stein, (2004) has shown that, if for some $a>0$, $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\;\left\{\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right\}^{2}\;{\rm d}z<\infty,$ (31)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\boldsymbol{s}),\boldsymbol{s}\in D$. Then (31) is the key tool to prove Theorem 1.
We first need the results in Lemma 2 hereunder. They follow from Zastavnyi, (2006), Chernih and Hubbert, (2014), and from standard properties of Fourier transforms and give asymptotic expansion of the spectral density of the $\mathcal{GH}$ model. Their proofs are thus omitted. For two given functions $g_{1}(x)$ and $g_{2}(x)$, with $g_{1}(x)\asymp g_{2}(x)$ we mean that there exist two constants $c$ and $C$ such that $0<c<C<\infty$ and $c|g_{2}(x)|\leq|g_{1}(x)|\leq C|g_{2}(x)|$ for each $x$.
@@
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be the function defined at Equation (11). Then, for $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $12(\beta+\gamma)\geq 6\delta+$ 1
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)=\sigma^{2}Ka^{d}\Big{[}c_{3}(za)^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
        $+\,c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\big{\{}\cos(za+c_{5})+\mathcal{O}(z^{-1})\big{\}}\Big{]}$,  for $z\to\infty$;
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta}$,  for $z\to\infty,$
where $c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)}$, $c_{4}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)}$ and $c_{5}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta-\gamma)$.
We are now ready to prove the Theorem. Let us start with the sufficient part of the assertion. From Lemma  2 (Point 2), we have:
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}\leq c_{2},\quad i=0,1.$
We then proceed by direct construction, and, using Lemma 2 (Points 1 and 2) and (31) we find that, as $z\to\infty$,
$\Big{|}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\Big{|}\leq\Big{|}\sigma_{1}^{2}K_{1}a_{1}^{d}\Big{[}c_{3}^{1}(za_{1})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$\displaystyle+\,c_{4}^{1}(za_{1})^{\lambda_{1}}\big{\{}\cos(za_{1}+c_{5}^{1})+\mathcal{O}(z^{-1})\big{\}}\Big{]}-\sigma_{0}^{2}K_{0}a_{0}^{d}\Big{[}c_{3}^{0}(za_{0})^{-2\delta}\big{\{}1+\mathcal{O}(z^{-2})\big{\}}$
$+\,c_{4}^{0}(za_{0})^{\lambda_{0}}\big{\{}\cos(za+c_{5}^{0})+\mathcal{O}(z^{-1})\big{\}}\Big{]}\Big{|}\frac{z^{2\delta}}{c_{1}}$
$\leq\frac{1}{c_{1}}\Big{|}\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\}+\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})$
$-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}+\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}\Big{|},$
where $c_{3}^{i}=\frac{2^{2\delta}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{\Gamma(\gamma_{i}-\delta)\Gamma(\beta_{i}-\delta)}$, $c_{4}^{i}=\frac{\pi^{-\frac{1}{2}}\Gamma(\beta_{i})\Gamma(\gamma_{i})}{2^{\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}}\Gamma(\delta)}$,$c_{5}^{i}=\frac{\pi}{2}(\frac{1}{2}+\delta-\beta_{i}-\gamma_{i})$ and $\lambda_{i}=\frac{1}{2}+\delta-\beta_{i}-\gamma_{i}$, $i=0,1$.
Let us now write
$A(z)$ $=\big{[}c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta)}K_{0}\big{]}\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big{[}c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z-c_{5}^{1})-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z-c_{5}^{0})\big{]}$
$D(z)$ $=\mathcal{O}(z^{-1})\{c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\}.$
Then, for some positive and finite $c$, Condition (31) can be written as
$\int_{c}^{\infty}z^{d-1}\bigg{(}\frac{\sigma_{1}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{1},\gamma_{1},a_{1}}(z)-\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}{\sigma_{0}^{2}\widehat{\mathcal{GH}}_{\delta,\beta_{0},\gamma_{0},a_{0}}(z)}\biggr{)}^{2}{\rm d}z<\infty,$ (32)
being equivalent to

### v2

@@ 6 Asymptotic properties of the maximum likelihood estimator of the Hypergeometric model
    $\sqrt{n}\big(\hat{\sigma}_{n}^{2}(\hat{a}_{n})/\hat{a}_{n}^{1+2\kappa}-\sigma_{0}^{2}/a_{0}^{1+2\kappa}\big)\stackrel{{\scriptstyle\mathcal{D}}}{{\longrightarrow}}\mathcal{N}\big(0,2(\sigma_{0}^{2}/a_{0}^{1+2\kappa})^{2}\big)$.
The condition $\mu\geq 4$ ensures the monotonicity property in Lemma 1 and is technical in nature.
Finally, we compare the two reparameterized models $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ and $\sigma^{2}\mathcal{H}_{\kappa,\mu,\alpha\mu,d}$. Under fixed-domain asymptotics and using (22), the associated microergodic parameters are
$\frac{\sigma^{2}}{\alpha^{2\kappa+1}}\quad\text{and}\quad\frac{\sigma^{2}B(\kappa,\mu,d)}{(\alpha\mu)^{2\kappa+1}},$
respectively. In the first case, the microergodic parameter does not depend on $\mu$, so both $\sigma^{2}/\alpha^{2\kappa+1}$ and $\mu$ are consistently estimable. From an inferential perspective, $\sigma^{2}\mathcal{H}_{\kappa,\mu,b(\alpha,\mu,\kappa,d),d}$ is therefore preferable. For large datasets, however, $\mu$ may be fixed to induce sparsity, in which case the microergodic parameter reduces to $\sigma^{2}/\alpha^{2\kappa+1}$ under both parameterizations.
@@ 7 A simulation study
We investigate the finite-sample performance of ML estimation for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$. We simulate 500 realizations of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ for $d=2$, $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. The process is generated via Cholesky decomposition on a regular grid with spacing $0.025$ over $0,1]^{2}$ (1681 locations), followed by independent perturbations of each coordinate by a uniform random variable on [0.01,0.01][-0.01,0.01].
Throughout the study, $\kappa$ and $\mu$ are assumed known. We focus on ML estimation of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing- and fixed-domain scenarios. Following Zhang and Zimmerman (2005); Kaufman and Shaby (2013), these regimes can be mimicked by fixing the sampling grid and varying the spatial dependence. Accordingly, we consider $a=0.1,0.25,0.5$, interpreting $a=0.1$ as increasing-domain and $a=0.5$ as fixed-domain behavior.
Table 1 reports the bias and standard deviation (in parentheses) of the 500 ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$, for $i=1,\ldots,500$.
@@
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.0150-0.0150 0.0219-0.0219 0.0151-0.0151 3.9684-3.9684 4.4120-4.4120 3.7659-3.7659 496.06-496.06 486.01-486.01 346.94-346.94
@@
Overall, the ML estimators of $\sigma^{2}$ and $a$ are approximately unbiased across all configurations. As expected, their standard deviations increase when moving from the increasing-domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$). For fixed $\kappa$, larger values of $\mu$ generally reduce variability.
Regarding the microergodic parameter, under the increasing-domain scenario ($a=0.1$) the estimator exhibits moderate bias and relatively high variance, particularly for larger $\kappa$. Under the fixed-domain scenario ($a=0.5$), both bias and variance decrease substantially, and the sampling distribution appears closer to Gaussian, in agreement with the asymptotic theory. Moreover, the variability of the microergodic estimator increases with the smoothness parameter $\kappa$.
Figure 4 displays boxplots of the ML estimates of $\sigma^{2}$, $a$, and $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (left) and $a=0.5$ (right), when $\kappa=0$. Under the increasing-domain scenario, the distributions of $\widehat{\sigma}^{2}$ and $\widehat{a}$ are approximately Gaussian. In the fixed-domain setting, the microergodic estimator becomes more concentrated around the true value and exhibits an approximately Gaussian distribution, as predicted by the theory.
@@ 8 A real data application
We consider data from WorldClim (www.worldclim.org), which provides high-resolution gridded climate surfaces for 1970–2000 (Fick and Hijmans, 2017) and can be accessed in R via geodata (Hijmans et al., 2024). We analyze September mean temperature over Madagascar. The full dataset consists of 725,330 grid-cell predictions obtained by interpolating weather-station data using thin-plate splines with geographic and satellite-derived covariates (Fick and Hijmans, 2017). To make ML estimation computationally feasible, we randomly sample 3,500 grid cells and use UTM-projected coordinates.
Following Li and Zhang (2011), the data are first detrended using spline regression to remove large-scale spatial trends. The residuals $y(\mathbf{s}_{i}),i=1,\ldots,3500$, are then treated as observations from a zero-mean RF $\{Y(\mathbf{s}):\mathbf{s}\in D\}$. Figure 5 displays the spatial map of the residuals and the empirical semivariogram. The plots suggest that a weakly stationary RF with a potentially heavy-tailed marginal distribution provides a reasonable modeling framework.
Refer to caption Refer to caption
Figure 5: Left: spatial map of residuals. Right: empirical semivariogram.
Let $Z(\mathbf{s})$ be a standard Gaussian RF with isotropic correlation function. We consider two specifications for $Y(\mathbf{s})$:
  • Gaussian RF: $Y(\mathbf{s})=\sigma Z(\mathbf{s})$;
  • Tukey-$h$ RF (Xu and Genton, 2017; Caamaño-Carrillo et al., 2024): $Y(\mathbf{s})=\sigma Z(\mathbf{s})\exp\!\left(\tfrac{hZ^{2}(\mathbf{s})}{2}\right)$, where $h\in[0,1/2)$ controls tail heaviness. Since the transformation is monotone, the Tukey-$h$ RF preserves the geometric properties of the underlying Gaussian field.
For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$, and the proposed ${\cal H}_{\kappa,\mu,\beta,2}$. Here $\kappa$ denotes smoothness, $\beta$ the scale parameter, and $\mu$ the shape parameter.
We adopt an increasing-domain framework and estimate all covariance parameters by ML. Standard errors are computed from the inverse Fisher information matrix. For the Tukey-$h$ RF, ML estimation is implemented via the Lambert-$W$ function (see equation 10 in Caamaño-Carrillo et al. (2024)).
Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values. Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model. Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$ achieve the best AIC values.
Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,\beta,2}$ correlation. The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$. Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
Gaussian RF Tukey-$h$ RF
@@
$\%$ 0 $96.45$ $96.23$ 0 $97.18$ $97.06$
TimeTime $7.22$ $0.04$ $0.04$ $6.26$ $0.03$ $0.03$
We assess predictive performance via repeated cross-validation. For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction. Table 2 reports the average RMSE and MAE. The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
@@
Figure 6: First row: normalized histogram versus estimated density and empirical semivariogram versus estimated semivariogram when using an estimated Gaussian RF with $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ correlation model (from left to right). Second row: the same graphics using a Tukey-$h$ RF.
@@ 9 Concluding Remarks
In this paper we study the Gauss hypergeometric covariance model, a flexible class introduced by Emery and Alegría (2022) that encompasses both the Matérn and Generalized Wendland families. We derive necessary and sufficient validity conditions and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics. To address this issue, we introduce a parsimonious compactly supported subclass selected through an optimality criterion based on maximum integral range. In this sense, the proposed hypergeometric model can be viewed as a structural refinement of the Generalized Wendland construction.
Although simple closed-form expressions are not generally available for nonnegative integer smoothness parameters, efficient numerical evaluation of the Gauss hypergeometric function is widely supported in modern scientific computing libraries. In particular, reliable implementations are available in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python. The GeoModels package (Bevilacqua et al., 2026), used in this paper for numerical experiments and applications, computes the proposed model via the SciPy implementation of the hypergeometric function (Virtanen et al., 2020).
@@ Acknowledgement
This work was funded by the National Agency for Research and Development of Chile, under grants ANID FONDECYT 1240308 (M. Bevilacqua), ANID project Data Observatory Foundation DO210001 (M. Bevilacqua), MATH-AMSUD1167 22-MATH-06 (AMSUD220041) (M. Bevilacqua), ANID FONDECYT 1250862 (C. Caamaño-Carrillo) and ANID PIA AFB230001 (X. Emery).
@@ A.1 Proof of Theorem 1
Let $P(\rho_{i})$, $i=0,1$, be two zero-mean Gaussian measures with isotropic covariance function $\rho_{i}$ and associated spectral density $\widehat{\rho}_{i}$, $i=0,1$, as defined through (1).
Using Skorokhod and Yadrenko (1973); Ibragimov and Rozanov (1978); Stein (2004), if for some $a>0$ the function $\widehat{\rho}_{0}(z)z^{a}$ is bounded away from 0 and $\infty$ as $z\to\infty$, and if for some finite and positive $c$,
$\int_{c}^{\infty}z^{d-1}\left(\frac{\widehat{\rho}_{1}(z)-\widehat{\rho}_{0}(z)}{\widehat{\rho}_{0}(z)}\right)^{2}\,{\rm d}z<\infty,$ (26)
then for any bounded subset $D\subset\mathbb{R}^{d}$, $P(\rho_{0})$ and $P(\rho_{1})$ are equivalent Gaussian measures on the paths of $Z(\mathbf{s})$, $\mathbf{s}\in D$. Condition (26) is the key tool.
We first need the following result.
@@
Let $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ be defined at (7). Assume $\delta>d/2$, $2(\beta-\delta)(\gamma-\delta)\geq\delta$ and $2(\beta+\gamma)\geq 6\delta+1$. Then:
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)$ $=\sigma^{2}Ka^{d}\Big[c_{3}(za)^{-2\delta}\{1+\mathcal{O}(z^{-2})\}$
    $\qquad+c_{4}(za)^{\frac{1}{2}+\delta-\beta-\gamma}\{\cos(za+c_{5})+\mathcal{O}(z^{-1})\}\Big],$
    as $z\to\infty$;
@@
    $\widehat{{\cal GH}}_{\delta,\beta,\gamma,a}(z)\asymp z^{-2\delta},\qquad z\to\infty.$
Here
$c_{3}=\frac{2^{2\delta}\Gamma(\beta)\Gamma(\gamma)}{\Gamma(\gamma-\delta)\Gamma(\beta-\delta)},\qquad c_{4}=\frac{\pi^{-1/2}\Gamma(\beta)\Gamma(\gamma)}{2^{\frac{1}{2}+\delta-\beta-\gamma}\Gamma(\delta)},$
$c_{5}=\frac{\pi}{2}\left(\frac{1}{2}+\delta-\beta-\gamma\right).$
We are now ready to prove the theorem.
From Lemma 2 (Point 2), there exist constants $0<c_{1}<c_{2}<\infty$ such that, for $z$ large,
$c_{1}\leq z^{2\delta}\widehat{{\cal GH}}_{\delta,\beta_{i},\gamma_{i},a_{i}}(z)\leq c_{2},\qquad i=0,1.$ (27)
In particular, $\widehat{\rho}_{0}(z)z^{2\delta}$ is bounded away from 0 and $\infty$ as $z\to\infty$.
Using Lemma 2 (Points 1 and 2), and collecting terms, one obtains for large $z$
$\left|\frac{\sigma_{1}^{2}\widehat{\varphi}_{1}(z)-\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}{\sigma_{0}^{2}\widehat{\varphi}_{0}(z)}\right|\leq\frac{1}{c_{1}}\left|A(z)+B(z)+D(z)\right|,$
where
$A(z)$ $=\big[c_{3}^{1}\sigma_{1}^{2}a_{1}^{d-2\delta}K_{1}-c_{3}^{0}\sigma_{0}^{2}a_{0}^{d-2\delta}K_{0}\big]\{1+\mathcal{O}(z^{-2})\},$
$B(z)$ $=\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}\cos(a_{1}z+c_{5}^{1})$
$\quad-c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\cos(a_{0}z+c_{5}^{0})\big],$
$D(z)$ $=\mathcal{O}(z^{-1})\big[c_{4}^{1}z^{\lambda_{1}+2\delta}\sigma_{1}^{2}a_{1}^{\lambda_{1}+d}K_{1}+c_{4}^{0}z^{\lambda_{0}+2\delta}\sigma_{0}^{2}a_{0}^{\lambda_{0}+d}K_{0}\big].$
Hence condition (26) becomes